# FedFACT: A Provable Framework for Controllable Group-Fairness Calibration in Federated Learning

### v1

@@ A.2 Group Fairness notations in FL
As noted in the main text, fairness at the level of each client’s dataset (local fairness) differs from fairness across the aggregate dataset of all clients (global fairness). Local fairness is defined with respect to each client’s individual data distribution $\mathbb{P}(X,Y,A\mid K)$, whereas global fairness is defined over the overall (aggregate) distribution $\mathbb{P}(X,Y,A)$. Motivated by approaches that employ group-specific confusion matrices for fairness [73, 55], we propose the decentralized group-specific confusion matrices $\mathbf{C}^{a,k},a\in\mathcal{A},k\in[N]$ to capture both global and local fairness across multiple data distributions within FL, with elements defined for $i,j\in[m]$ as $\mathbf{C}_{i,j}^{a,k}(h):=\mathbb{P}(Y=i,\widehat{Y}=j\mid A=a,K=k)$.
@@
where $\mathbf{D}^{a,k}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$, and the $y$-th column elements of $\mathbf{D}^{a,k}_{a^{\prime},y}$ are $\mathbb{P}(K=k\mid A=a^{\prime})\mathbb{I}[a=a^{\prime}]-\mathbb{P}(A=a,K=k)$ with all other elements set to 00.
@@
$\mathscr{D}_{DP}^{l,k}=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\left|\mathbb{P}(\widehat{Y}=y\mid A=a^{\prime},K=k)-\mathbb{P}(\widehat{Y}=y\mid K=k)\right|,$
@@
$\mathscr{D}_{DP}^{l,k}=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\left|\sum_{a\in\mathcal{A}}\sum_{i\in[m]}\big{(}\mathbb{I}[a=a^{\prime}]-p_{a\mid k}\big{)}\mathbf{C}^{a,k}_{i,y}(h_{k})\right|=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\left|\sum_{a\in\mathcal{A}}\langle\mathbf{D}^{a,k}_{a^{\prime},y},\mathbf{C}^{a,k}(h_{k})\rangle\right|,$
where $\mathbf{D}^{a,k}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$, and the $y$-th column elements of $\mathbf{D}^{a,k}_{a^{\prime},y}$ are $\mathbb{I}[a=a^{\prime}]-\mathbb{P}(A=a\mid K=k)$ with all other elements set to 00.
@@
where $\mathbf{D}^{a,k}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$, and the entry in the $y$-th row and $y$-th column is $\frac{p_{a^{\prime},k}}{p_{a^{\prime},y}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y}}$ with all other elements set to 00.
@@
$\mathscr{D}_{EOP}^{l,k}=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\left|\mathbb{P}(\widehat{Y}=y\mid A=a^{\prime},Y=y,K=k)-\mathbb{P}(\widehat{Y}=y\mid Y=y,K=k)\right|,$
@@
$\mathscr{D}_{EOP}^{l,k}=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\left|\sum_{a\in\mathcal{A}}\bigg{(}\frac{p_{a^{\prime},k}}{p_{a^{\prime},y,k}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y,k}}\bigg{)}\mathbf{C}^{a,k}_{i,y}(h_{k})\right|=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\left|\sum_{a\in\mathcal{A}}\langle\mathbf{D}^{a,k}_{a^{\prime},y},\mathbf{C}^{a,k}(h_{k})\rangle\right|,$
where $\mathbf{D}^{{a,k}}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$, and the entry in the $y$-th row and $y$-th column is $\frac{p_{a^{\prime},k}}{p_{a^{\prime},y,k}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y,k}}$ with all other elements set to 00.
@@
For EO criterion, the global EO fairness metric is defined as
@@
where the entry in the $y$-th row and $y$-th column is $\frac{a^{\prime},k}{p_{a^{\prime},y}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y}}$ with all other elements set to 00 for $\mathbf{D}^{a,k,0}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$, and the entry in the $y$-th column is $\frac{p_{a^{\prime},k}}{\sum_{y_{j}\neq y}p_{a^{\prime},y_{j}}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{\sum_{y_{j}\neq y}p_{y_{j}}}$ except for the $y$-th row with all other elements set to 00 for $\mathbf{D}^{a,k,1}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$.
The local EO fairness metric for $k$-th client is defined as
$\mathscr{D}_{EO}^{l,k}=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\frac{1}{2}\bigg{(}$ $\left|\mathbb{P}(\widehat{Y}=y\mid Y=y,A=a^{\prime},K=k)-\mathbb{P}(\widehat{Y}=y\mid Y=y,K=k)\right|$
@@
where the entry in the $y$-th row and $y$-th column is $\frac{a^{\prime},k}{p_{a^{\prime},y,k}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y,k}}$ with all other elements set to 00 for $\mathbf{D}^{a,k,0}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$, and the entry in the $y$-th column is $\frac{p_{a^{\prime},k}}{\sum_{y_{j}\neq y}p_{a^{\prime},y_{j},k}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{\sum_{y_{j}\neq y}p_{y_{j},k}}$ except for the $y$-th row with all other elements set to 00 for $\mathbf{D}^{a,k,1}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$.
@@ B.1 Proof of Proposition 1
This section provides the proof of Proposition 1. The proof is primarily inspired by the characterization of the Bayes-optimal fair classifier in the centralized fair machine learning literature (e.g. Theorem 3.1 of [73], Proposition 10 of [55]).
@@
where $\mathcal{C}^{|\mathcal{A}|\times N}$ be the product space of all confusion matrices $\mathbf{C}^{a,k}$ corresponding to sensitive group $a\in\mathcal{A}$ and $k\in[N]$ associated with a given instance $\mathbf{h}\in\mathcal{H}^{N}$ of the problem. It is clear that the performance metric $\mathcal{R}$ and fairness metrics $\mathscr{D}^{g},\mathscr{D}^{l,k},k\in[N]$ are continuous and bounded to $\mathbf{C}^{|\mathcal{A}|\times N}(\mathbf{h}):=\{\mathbf{C}^{a,k}(h_{k})\}_{a\in\mathcal{A},k\in[N]}$.
@@
Deterministic classifiers. It can be seen that, for any linear objective $\phi_{\mathbf{L}}(\mathbf{C}^{|\mathcal{A}|\times N}(\mathbf{h}))=\sum_{a\in\mathcal{A}}\sum_{k\in[N]}\langle\mathbf{L}^{a,k},\mathbf{C}^{a,k}(\mathbf{h}_{k})\rangle$, there is a deterministic classifiers $\mathbf{h}^{*}=({h}^{*}_{1},\cdots,{h}^{*}_{N})$ that is optimal for $\phi_{\mathbf{L}}$ (see proof in B.2). By the supporting-hyperplane theorem [8] for compact convex sets, for each point $\mathbf{C}_{b}=\{\mathbf{C}^{a,k}_{b}\}_{a\in\mathcal{A},k\in[N]}\in\partial\mathcal{C}^{|\mathcal{A}|\times N}$, there exists a nonzero collection of matrices $\mathbf{L}_{b}=\{\mathbf{L}^{a,k}_{b}\}_{a\in\mathcal{A},\,k\in[N]}$ constitutes a hyperplane, such that for every $\mathbf{C}=\{\mathbf{C}^{a,k}\}\in\mathcal{C}^{|\mathcal{A}|\times N}$ we have $\sum_{a\in\mathcal{A}}\sum_{k=1}^{N}\langle\mathbf{L}^{a,k}_{b},\mathbf{C}^{a,k}_{b}\rangle\;\leq\;\sum_{a\in\mathcal{A}}\sum_{k=1}^{N}\langle\mathbf{L}^{a,k}_{b},\mathbf{C}^{a,k}\rangle\,$ which is precisely the desired supporting-hyperplane condition at $\mathbf{C}_{b}$. In other words, we arrive at the conclusion that each boundary point of $\mathcal{C}^{|\mathcal{A}|\times N}$ can be achieved by deterministic classifiers $\mathbf{h}^{\prime}=(h^{\prime}_{1},\cdots,h^{\prime}_{N})$.
@@
Combination of deterministic classifiers. Since $\mathcal{C}^{|\mathcal{A}|\times N}$ is compact and convex, we know that its extreme points fall in its boundary. By the Krein-Milman theorem [56], we have that $\mathcal{C}^{|\mathcal{A}|\times N}$ is equal to the convex hull of its extreme points. We further have from Caratheodory’s theorem [8] that any $\mathbf{C}\in\mathcal{C}^{|\mathcal{A}|\times N}$ can be expressed as a convex combination of $d_{k}=|\mathcal{A}|Nm^{2}$ points in the extreme point set, where each extreme point can be characterized by deterministic classifiers. Hence, we have proved that the optimal solution $\mathbf{h}$ can be represented by the convex combination of deterministic classifiers. $\square$
@@
Discussion on feasibility. The only condition for the above theorem to hold is that the feasible set is non-empty, which is clearly satisfied by the mentioned fairness constraints, DP, EOP, and EO. For these fairness criteria, the classifier that always predicts a single, fixed label $y^{\prime}$ trivially meets $\xi^{g}=0,\xi^{l,k}=0,k\in[N]$, and hence satisfies the fairness constraints.
@@
The number of deterministic classifiers. As for the number of deterministic classifiers required, the parameter $d^{k}$ in the proof scales with the number of nonzero entries in the linear performance and fairness constraints [55]. Since each matrix $\mathbf{D}^{a,k}$ in our fairness formulation is zero except for one column, we in fact need far fewer than $|\mathcal{A}|Nm^{2}$ classifiers. Moreover, under the continuity assumption 1, this number can be reduced even further [73].
@@ B.2 Proof of Proposition 2
$\quad+\sum_{k=1}^{N}\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu^{(1)}_{k,u_{l,k}}-\mu^{(2)}_{k,u_{l,k}})\sum_{a\in\mathcal{A}}\langle\mathbf{D}^{a,k}_{u_{l,k}},\mathbf{C}^{a,k}(h_{k})\rangle-\sum_{{u_{g}}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}+\lambda^{(2)}_{u_{g}})\xi^{g}$
$\quad-\sum_{k\in[N]}\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu^{(1)}_{k,u_{l,k}}+\mu^{(2)}_{k,u_{l,k}})\xi^{l,k}$
$=\sum_{k=1}^{N}\sum_{a\in\mathcal{A}}\Bigg{\langle}p_{a,k}(\mathbf{1}-\mathbf{I})+\sum_{{u_{g}}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}-\lambda^{(2)}_{u_{g}})\mathbf{D}^{a,k}_{u_{g}}+\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu^{(1)}_{k,u_{l,k}}-\mu^{(2)}_{k,u_{l,k}})\mathbf{D}^{a,k}_{u_{l,k}},\mathbf{C}^{a,k}(h_{k})\Bigg{\rangle}$
$\quad-\sum_{{u_{g}}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}+\lambda^{(2)}_{u_{g}})\xi^{g}-\sum_{k\in[N]}\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu^{(1)}_{k,u_{l,k}}+\mu^{(2)}_{k,u_{l,k}})\xi^{l,k}.$
@@
where $\mathbf{M}^{\lambda,\mu}(a,k):=\mathbf{I}-\frac{1}{p_{a,k}}\Big{[}\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}-\lambda^{(2)}_{u_{g}})\mathbf{D}^{a,k}_{u_{g}}-\sum_{u_{l,k}\in\mathcal{U}_{l,k}}(\mu^{(1)}_{k,u_{l,k}}-\mu^{(2)}_{k,u_{l,k}})\mathbf{D}^{a,k}_{u_{l,k}}\Big{]}$.

### v2

@@ A.2 Group Fairness notations in FL
As noted in the main text, fairness at the level of each client’s dataset (local fairness) differs from fairness across the aggregate dataset of all clients (global fairness). Local fairness is defined with respect to each client’s individual data distribution $\mathbb{P}(X,Y,A\mid K)$, whereas global fairness is defined over the overall (aggregate) distribution $\mathbb{P}(X,Y,A)$. Motivated by approaches that employ group-specific confusion matrices for fairness [overlap, coonfusion-jmlr-complex-constraints], we propose the decentralized group-specific confusion matrices $\mathbf{C}^{a,k},a\in\mathcal{A},k\in[N]$ to capture both global and local fairness across multiple data distributions within FL, with elements defined for $i,j\in[m]$ as $\mathbf{C}_{i,j}^{a,k}(h):=\mathbb{P}(Y=i,\widehat{Y}=j\mid A=a,K=k)$.
@@
where $\mathbf{D}^{a,k}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$, and the $y$-th column elements of $\mathbf{D}^{a,k}_{a^{\prime},y}$ are $\mathbb{P}(K=k\mid A=a^{\prime})\mathbb{I}[a=a^{\prime}]-\mathbb{P}(A=a,K=k)$ with all other elements set to 0.
@@
$\mathscr{D}_{DP}^{k}=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\left|\mathbb{P}(\widehat{Y}=y\mid A=a^{\prime},K=k)-\mathbb{P}(\widehat{Y}=y\mid K=k)\right|,$
@@
$\mathscr{D}_{DP}^{k}=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\left|\sum_{a\in\mathcal{A}}\sum_{i\in[m]}\big(\mathbb{I}[a=a^{\prime}]-p_{a\mid k}\big)\mathbf{C}^{a,k}_{i,y}(h_{k})\right|=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\left|\sum_{a\in\mathcal{A}}\langle\mathbf{D}^{a,k}_{a^{\prime},y},\mathbf{C}^{a,k}(h_{k})\rangle\right|,$
where $\mathbf{D}^{a,k}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$, and the $y$-th column elements of $\mathbf{D}^{a,k}_{a^{\prime},y}$ are $\mathbb{I}[a=a^{\prime}]-\mathbb{P}(A=a\mid K=k)$ with all other elements set to 0.
@@
where $\mathbf{D}^{a,k}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$, and the entry in the $y$-th row and $y$-th column is $\frac{p_{a^{\prime},k}}{p_{a^{\prime},y}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y}}$ with all other elements set to 0.
@@
$\mathscr{D}_{EOP}^{k}=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\left|\mathbb{P}(\widehat{Y}=y\mid A=a^{\prime},Y=y,K=k)-\mathbb{P}(\widehat{Y}=y\mid Y=y,K=k)\right|,$
@@
$\mathscr{D}_{EOP}^{k}=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\left|\sum_{a\in\mathcal{A}}\bigg(\frac{p_{a^{\prime},k}}{p_{a^{\prime},y,k}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y,k}}\bigg)\mathbf{C}^{a,k}_{i,y}(h_{k})\right|=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\left|\sum_{a\in\mathcal{A}}\langle\mathbf{D}^{a,k}_{a^{\prime},y},\mathbf{C}^{a,k}(h_{k})\rangle\right|,$
where $\mathbf{D}^{{a,k}}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$, and the entry in the $y$-th row and $y$-th column is $\frac{p_{a^{\prime},k}}{p_{a^{\prime},y,k}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y,k}}$ with all other elements set to 0.
@@
For EOP criterion, the global EOP fairness metric is defined as
@@
where the entry in the $y$-th row and $y$-th column is $\frac{a^{\prime},k}{p_{a^{\prime},y}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y}}$ with all other elements set to 0 for $\mathbf{D}^{a,k,0}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$, and the entry in the $y$-th column is $\frac{p_{a^{\prime},k}}{\sum_{y_{j}\neq y}p_{a^{\prime},y_{j}}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{\sum_{y_{j}\neq y}p_{y_{j}}}$ except for the $y$-th row with all other elements set to 0 for $\mathbf{D}^{a,k,1}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$.
The local EOP fairness metric for $k$-th client is defined as
$\mathscr{D}_{EO}^{k}=\max_{y\in[m]}\max_{a^{\prime}\in\mathcal{A}}\frac{1}{2}\bigg($ $\left|\mathbb{P}(\widehat{Y}=y\mid Y=y,A=a^{\prime},K=k)-\mathbb{P}(\widehat{Y}=y\mid Y=y,K=k)\right|$
@@
where the entry in the $y$-th row and $y$-th column is $\frac{a^{\prime},k}{p_{a^{\prime},y,k}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{p_{y,k}}$ with all other elements set to 0 for $\mathbf{D}^{a,k,0}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$, and the entry in the $y$-th column is $\frac{p_{a^{\prime},k}}{\sum_{y_{j}\neq y}p_{a^{\prime},y_{j},k}}\mathbb{I}[a=a^{\prime}]-\frac{p_{a,k}}{\sum_{y_{j}\neq y}p_{y_{j},k}}$ except for the $y$-th row with all other elements set to 0 for $\mathbf{D}^{a,k,1}_{a^{\prime},y}\in\mathbb{R}^{m\times m}$.
A.3 Other Fairness Notations
Fairness notions formulated as ratios metrics can be converted into linear constraints under certain conditions, and our framework is well suited to enforce these fairness constraints in federated learning environments. Specifically, the ratios metric or constraints, which are formulated as $\left|\frac{\sum_{a}\langle\mathbf{D}^{a},\mathbf{C}^{a}(h)\rangle}{\sum_{a}\langle\mathbf{G}^{a},\mathbf{C}^{a}(h)\rangle}\right|\leq\xi$ with constant metrix $\mathbf{D}^{a},\mathbf{G}^{a}$ and group-specific confusion matrix $\mathbf{C}^{a}(h)$ depend on classifier $h$, can certainly be transformed into multiple linear constraints if the sign of $\sum_{a}\langle\mathbf{G}^{a},\mathbf{C}^{a}(h)\rangle$ is unchanged for any $h$. When the denominator’s sign is uncertain, the feasible domain of $\mathbf{C}^{a}(h)$ is non‑convex, precluding its expression via linear constraints. In fact, since each entry of $\mathbf{C}^{a}(h)$ lies in [0,1], whenever the entries of $\mathbf{G}^{a}$ are sign‑consistent, the corresponding ratio constraint admits a linear‑constraint representation. For example, the Calibration within Groups (CG) which was proposed in [Fairness-in-Criminal] and further explored in [FACT-confusion], is a fairness metric in binary classification and can be formulated as $\frac{FN^{a}}{FN^{a}+TN^{a}}=v_{0};\frac{TP^{a}}{TP^{a}+FP^{a}}=v_{1}$, where $TP^{a},FP^{a},FN^{a},TN^{a}$ are derived from binary group-specific confusion matrix $\mathbf{C}^{a}$ and $0\leq v_{0}<v_{1}\leq 1$ and have no implicit dependence on any entries of the fairness-confusion tensor. Because this fairness criterion appears as a ratio metric and every element of corresponding $\mathbf{G}^{a}$ is non‑negative, it admits a linear‑constraint representation and can be realized in our proposed distributed framework. Moreover, the ratio metrics presented in [beyond-alghamdi2022beyond] also can be formulated into multiple linear constraints based on the above analysis.
@@ B.1 Proof of Proposition 1
This section provides the proof of Proposition 1. The proof is primarily inspired by the characterization of the Bayes-optimal fair classifier in the centralized fair machine learning literature (e.g. Theorem 3.1 of [overlap], Proposition 10 of [coonfusion-jmlr-complex-constraints]).
@@
where $\mathcal{C}^{|\mathcal{A}|\times N}$ be the product space of all confusion matrices $\mathbf{C}^{a,k}$ corresponding to sensitive group $a\in\mathcal{A}$ and $k\in[N]$ associated with a given instance $\mathbf{h}\in\mathcal{H}^{N}$ of the problem. It is clear that the performance metric $\mathcal{R}$ and fairness metrics $\mathscr{D}^{g},\mathscr{D}^{k},k\in[N]$ are continuous and bounded to $\mathbf{C}^{|\mathcal{A}|\times N}(\mathbf{h}):=\{\mathbf{C}^{a,k}(h_{k})\}_{a\in\mathcal{A},k\in[N]}$.
@@
Deterministic classifiers. It can be seen that, for any linear objective $\phi_{\mathbf{L}}(\mathbf{C}^{|\mathcal{A}|\times N}(\mathbf{h}))=\sum_{a\in\mathcal{A}}\sum_{k\in[N]}\langle\mathbf{L}^{a,k},\mathbf{C}^{a,k}(\mathbf{h}_{k})\rangle$, there is a deterministic classifiers $\mathbf{h}^{*}=({h}^{*}_{1},\cdots,{h}^{*}_{N})$ that is optimal for $\phi_{\mathbf{L}}$ (see proof in B.2). By the supporting‐hyperplane theorem [Convex-optimization-textbook-boyd2004convex] for compact convex sets, for each point $\mathbf{C}_{b}=\{\mathbf{C}^{a,k}_{b}\}_{a\in\mathcal{A},k\in[N]}\in\partial\mathcal{C}^{|\mathcal{A}|\times N}$, there exists a nonzero collection of matrices $\mathbf{L}_{b}=\{\mathbf{L}^{a,k}_{b}\}_{a\in\mathcal{A},\,k\in[N]}$ constitutes a hyperplane, such that for every $\mathbf{C}=\{\mathbf{C}^{a,k}\}\in\mathcal{C}^{|\mathcal{A}|\times N}$ we have $\sum_{a\in\mathcal{A}}\sum_{k=1}^{N}\langle\mathbf{L}^{a,k}_{b},\mathbf{C}^{a,k}_{b}\rangle\;\leq\;\sum_{a\in\mathcal{A}}\sum_{k=1}^{N}\langle\mathbf{L}^{a,k}_{b},\mathbf{C}^{a,k}\rangle\,$ which is precisely the desired supporting‐hyperplane condition at $\mathbf{C}_{b}$. In other words, we arrive at the conclusion that each boundary point of $\mathcal{C}^{|\mathcal{A}|\times N}$ can be achieved by deterministic classifiers $\mathbf{h}^{\prime}=(h^{\prime}_{1},\cdots,h^{\prime}_{N})$.
@@
Combination of deterministic classifiers. Since $\mathcal{C}^{|\mathcal{A}|\times N}$ is compact and convex, we know that its extreme points fall in its boundary. By the Krein-Milman theorem [Topological-vector-textbook-narici2010topological], we have that $\mathcal{C}^{|\mathcal{A}|\times N}$ is equal to the convex hull of its extreme points. We further have from Caratheodory’s theorem [Convex-optimization-textbook-boyd2004convex] that any $\mathbf{C}\in\mathcal{C}^{|\mathcal{A}|\times N}$ can be expressed as a convex combination of $d_{k}=|\mathcal{A}|Nm^{2}$ points in the extreme point set, where each extreme point can be characterized by deterministic classifiers. Hence, we have proved that the optimal solution $\mathbf{h}$ can be represented by the convex combination of deterministic classifiers. $\square$
@@
Discussion on feasibility. The only condition for the above theorem to hold is that the feasible set is non‐empty, which is clearly satisfied by the mentioned fairness constraints, DP, EOP, and EO. For these fairness criteria, the classifier that always predicts a single, fixed label $y^{\prime}$ trivially meets $\xi^{g}=0,\xi^{k}=0,k\in[N]$, and hence satisfies the fairness constraints.
@@
The number of deterministic classifiers. As for the number of deterministic classifiers required, the parameter $d^{k}$ in the proof scales with the number of nonzero entries in the linear performance and fairness constraints [coonfusion-jmlr-complex-constraints]. Since each matrix $\mathbf{D}^{a,k}$ in our fairness formulation is zero except for one column, we in fact need far fewer than $|\mathcal{A}|Nm^{2}$ classifiers. Moreover, under the continuity assumption 1, this number can be reduced even further [overlap].
@@ B.2 Proof of Proposition 2
$\quad+\sum_{k=1}^{N}\sum_{u_{k}\in\mathcal{U}_{k}}(\mu^{(1)}_{k,u_{k}}-\mu^{(2)}_{k,u_{k}})\sum_{a\in\mathcal{A}}\langle\mathbf{D}^{a,k}_{u_{k}},\mathbf{C}^{a,k}(h_{k})\rangle-\sum_{{u_{g}}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}+\lambda^{(2)}_{u_{g}})\xi^{g}$
$\quad-\sum_{k\in[N]}\sum_{u_{k}\in\mathcal{U}_{k}}(\mu^{(1)}_{k,u_{k}}+\mu^{(2)}_{k,u_{k}})\xi^{k}$
$=\sum_{k=1}^{N}\sum_{a\in\mathcal{A}}\Bigg\langle p_{a,k}(\mathbf{1}-\mathbf{I})+\sum_{{u_{g}}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}-\lambda^{(2)}_{u_{g}})\mathbf{D}^{a,k}_{u_{g}}+\sum_{u_{k}\in\mathcal{U}_{k}}(\mu^{(1)}_{k,u_{k}}-\mu^{(2)}_{k,u_{k}})\mathbf{D}^{a,k}_{u_{k}},\mathbf{C}^{a,k}(h_{k})\Bigg\rangle$
$\quad-\sum_{{u_{g}}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}+\lambda^{(2)}_{u_{g}})\xi^{g}-\sum_{k\in[N]}\sum_{u_{k}\in\mathcal{U}_{k}}(\mu^{(1)}_{k,u_{k}}+\mu^{(2)}_{k,u_{k}})\xi^{k}.$
@@
where $\mathbf{M}^{\lambda,\mu}(a,k):=\mathbf{I}-\frac{1}{p_{a,k}}\Big[\sum_{u_{g}\in\mathcal{U}_{g}}(\lambda^{(1)}_{u_{g}}-\lambda^{(2)}_{u_{g}})\mathbf{D}^{a,k}_{u_{g}}-\sum_{u_{k}\in\mathcal{U}_{k}}(\mu^{(1)}_{k,u_{k}}-\mu^{(2)}_{k,u_{k}})\mathbf{D}^{a,k}_{u_{k}}\Big]$.